# Estimation of AMOC transition probabilities using a machine learning based rare-event algorithm

### v1

@@ 6 Discussion
This is to our knowledge the first study to couple TAMS to a machine learning technique that estimates the committor function on-the-fly. On the one hand, Lucente et al. (2022) had already coupled Adaptive Multilevel Splitting (an algorithm almost similar to TAMS) to another machine learning technique, the Analogue Markov Chain, but the committor had to be estimated beforehand, thus creating an additional computational cost. On the other hand, direct committor function computation using machine learning has already been performed on complex climate data (Miloshevich et al., 2023b), but only with the objective of short-term probabilistic forecast.
@@
We have coupled TAMS to RC to show that the committor function could be learned at no additional cost than what is necessary to compute transition probabilities. We compared our results to a combination of TAMS experiments with a physics-informed score function and showed that the transition probabilities, the mean first-passage times and the transition paths were consistent. In the case of the F-transition, TAMS-$R$ also yields on average a similar variance in the results as TAMS-$S$, except for very low noise. Regarding the S-transition, both methods remain consistent for transition probabilities within $1000$ years down to $10^{-6}$. Furthermore, TAMS-$R$ consistently requires much fewer time steps to converge than TAMS-$S$, suggesting that in this case RCs can learn a more efficient score function than $\Phi_{S}^{S}$. Finally, we showed that analytical formulas can be derived for the learned committor are consistent in the case of the F-transition with the physics-informed score function.
@@
The next step is to study AMOC tipping in models of much higher dimensions, such as Earth Models of Intermediate Complexity (EMICs) or even Global Climate Models (GCMs). Another rare-event algorithm (Moral and Garnier, 2005; Tailleur and Kurchan, 2006; Lecomte and Tailleur, 2007) has already been applied to EMICs (Ragone et al., 2018; Ragone and Bouchet, 2021; Wouters et al., 2023) and to AMOC tipping in particular (Cini et al., 2023). It relies on other selection/deletion rules than a score function, but makes it more difficult to compute transition probabilities. Another interesting data-driven approach of AMOC tipping is the application of the edge-tracking algorithm by Mehling et al. (2024). It aims at finding the critical state through which an AMOC tipping would occur, using a limited number of simulations, and would thus be tractable in higher dimension. This approach provides insight into the dynamics of this system but not on quantities such as the probability of tipping or transition times.
@@
Extending our method to complex climate models raises the question of its scaling with the dimension of the model. For TAMS-$R$, this question is twofold: the feasibility of simulating enough trajectories on the one hand, and the scaling of RC itself, on the other. Regarding the first point, TAMS is not more computationally demanding than any other rare-event algorithm so their recent successes in EMICs (Cini et al., 2023) are promising for TAMS. When it comes to GCMs (van Westen and Dijkstra, 2023), running TAMS does not seem feasible yet.
@@
However, the current implementation of the RC method poses a real scaling problem. Using monomials as a family of basis functions results in a feature vector whose size grows exponentially with the number of variables. We can only consider a very limited number of variables, but it raises in turn the question whether such simplification is sufficient to describe the dynamics of AMOC tipping in a GCM. Miloshevich et al. (2023a) recently showed that Convolutional Neural Networks (CNN) are suitable to learn the committor from maps of climate variables by reducing them into a lower-dimension vector. However, the problem remains that CNNs require a large amount of data to be trained. More generally, this raises the problem of the choice of basis functions for the RC. The same issue arises for another committor estimation method, the Dynamical Galerkin Approximation (Thiede et al., 2019; Finkel et al., 2021; Jacques-Dumas et al., 2023). It is interesting to note that Strahan et al. (2023) recently used a Feedforward Neural Network to solve the Feynman-Kac formula for the committor. This approach only requires short trajectories such as those that can be obtained with TAMS but the issue of the amount of data and the training of the neural network remains.
@@
The interpretability of the RC could also prove useful on the question of dimensionality. For example, we showed in Section 5 that $\mathbf{W}_{\mathrm{out}}$ contains forty basis terms but only about ten play a significant role in the final score function. If these can be detected early enough in the learning process, dropping them would greatly speed up the computation. More generally, neural networks are often considered black boxes and difficult to interpret. Here, it is possible, by estimating analytical formulas, to highlight important variables for the system dynamics and their relationship. It must be used with caution, but can serve as a useful starting point to study the dynamics of the system or to focus observations or experiments. The efficiency of TAMS is considered to rely on sufficient insight into the model’s dynamics (Lestang et al., 2018), but we show here that an efficient TAMS can also be used to obtain such useful insight.
@@
Finally, the coupling between TAMS and RC must be refined, especially the initialization and training process of the latter. In complex models, transition probabilities may only be computed for a single model setup, which makes transfer learning unusable. The RC could then be initialized from a linear score function that makes it easy to generate surrogate data. Since we might not be able to make a better guess of a prescribed score function, this could be a suitable starting point for the learning process. Another important issue is the memory of the RC. The current transfer learning process probably may retain too much information corresponding to very different system parameters, since we did not ”clean” its memory during the computations. This problem does not only arise with transfer learning, and we would also encounter it if the initialization of the RC is very far from the actual committor. A possible solution could be to reinitialize the RC at certain milestones, once a certain number of transitions have been reached, thus using ”better” data.
@@ Acknowledgements.
    Generate $N$ trajectories $\left(\mathbf{X}^{(1)},\dots,\mathbf{X}^{(N)}\right)$ starting in set $A$ until $T_{\mathrm{max}}$ or until they reach set $B$. Repeat steps $3-9$ until $N-n_{c}$ trajectories have reached $B$ (or $N$ if $n_{c}=1$).
@@
    For trajectory $i$, the maximum value of the score function is called $Q_{i}$.
@@
    Let $I_{k}=\{i\in[1,N]\ |\ Q_{i}\in\{Q^{*}_{1},...,Q^{*}_{n_{c}}\}\}$, where $\{Q^{*}_{j}\}_{j\in[1,n_{c}]}$ are the $n_{c}$ lowest values of $Q_{i}$. Since the system is discrete, $I$ may contain more than $n_{c}$ elements.
@@
    Randomly select one of the remaining trajectories, say $\mathbf{X}^{(r)}$ (such that $r\notin I_{k}$). We call $\tau$ the first time so that $\Phi(\mathbf{X}^{(r)}(\tau))=Q_{i}$.
@@
We call $N_{B}$ the final number of trajectories that contain a transition, and the total number of iterations is called $K$. The estimated transition probability from $A$ to $B$ before $T_{\mathrm{max}}$ is given by the equation (Rolland and Simonnet, 2015; Rolland, 2022):
$\hat{q}=\frac{N_{B}}{N}w_{K}=\frac{N_{B}}{N}\prod_{k=1}^{K}w_{k}=\frac{N_{B}}{N}\prod_{k=0}^{K-1}\left(1-\frac{\#I_{k}}{N}\right)$ (9)
The optimal variance, obtained when the score function is the committor function (see Appendix B), reads:
$\sigma^{2}_{\mathrm{id}}=\frac{<\hat{q}>^{2}}{N}\left(<K>\frac{n_{c}}{N-n_{c}}+\frac{N-<N_{B}>}{<N_{B}>}\right)$ (10)
where <>$\cdot$>< ⋅ > represents an average over the $M$ runs of TAMS.
@@
Consider a stochastic dynamical system of dimension $d$ and two subsets $\mathcal{E}$ and $\mathcal{F}$ of its phase space. If the drift of the system is called $a$ and its diffusivity $D$, then the adjoint $\mathcal{L}$ of the Fokker-Planck operator is the infinitesimal generator of the stochastic process and defined as:
$\mathcal{L}=\frac{\partial}{\partial t}(\cdot)-\sum_{i=1}^{n}a_{i}(\mathbf{x})\frac{\partial}{\partial\mathbf{x}_{i}}(\cdot)-\sum_{i,j=1}^{n}D_{ij}(\mathbf{x})\frac{\partial}{\partial\mathbf{x}_{i}\partial\mathbf{x}_{j}}(\cdot),$ (11)
Formally, the committor function, as solution to the backward Kolmogorov equation, solves the Dirichlet problem (Lucente et al., 2022):
$\mathcal{L}\Phi(t,\mathbf{x})$ $=0\ \forall(t,\mathbf{x})\in\mathbb{R}\times(\mathcal{E}\cup\mathcal{F})^{c}$ (12)
$\Phi(t,\mathbf{x})$ $=1\ \forall(t,\mathbf{x})\in\mathbb{R}\times\mathcal{E}$
$\Phi(t,\mathbf{x})$ $=0\ \forall(t,\mathbf{x})\in\mathbb{R}\times\mathcal{F}$
The committor function $\Phi(t,\mathbf{x})$, as defined by the system (12) and its boundary conditions, also corresponds to the probability that a trajectory $\mathbf{X}(t)$ starting on $\mathbf{x}$ reaches the set $\mathcal{E}$ before $\mathcal{F}$. More precisely, for any set $\mathcal{G}\in\mathbb{R}^{n}$, the first passage time in $\mathcal{G}$ is defined as:
$\tau_{\mathcal{G}}=\min\{t\ |\ \mathbf{x}(t)\in\mathcal{G}\}$ (13)
The committor is defined as the probability that the first-passage time in $\mathcal{E}$ is lower than that in $\mathcal{F}$:
$\Phi(t,\mathbf{x})=\mathbb{P}(\tau_{\mathcal{E}}<\tau_{\mathcal{F}}\ |\ \mathbf{X}(0)=\mathbf{x})$ (14)
[…]
Reservoir Computing (Jaeger, 2001; Lukoševičius and Jaeger, 2009) is a kind of neural network that is designed to better represent dynamical systems and temporal processes than classical feedforward neural networks, without using gradient descent for training.
Reservoir Computing aims at transforming a nonlinear problem in the original phase space into a linear problem by linearly embedding input data into a larger-dimensional space. Then a linear regression is performed in this feature space. The main idea is that the matrix representing this space is random and fixed; only the linear regression parameters are learned during training. Although this neural network acts as a universal approximator (Gonon and Ortega, 2020), its use encounters certain problems, such as the randomness of the reservoir: different realizations of a random network can greatly change the performance of the reservoir computer.
@@
To solve this issue, Gauthier et al. (2021) developed another universal approximator (Gonon and Ortega, 2020; Hart et al., 2021) that they call “Next-Generation” Reservoir Computing (here called RC). The difference is that the input data is now embedded into a larger-dimensional space using a family of nonlinear basis functions instead of a random matrix. We provide below only the important steps to build an RC for our problem. Details of the architecture can be found in Gauthier et al. (2021) and in Jacques-Dumas et al. (2023) for the setup and choice of parameters.
@@
At each time step, the RC is given a vector $\mathbf{x}\in\mathbb{R}^{d}$ that contains here only $S_{n}$, $S_{ts}$, $D$ and time $t$. The family of non-linear functions consists of products of power $p$ of these variables. The resulting feature vector $\mathbf{\tilde{x}}\in\mathbb{R}^{M}$, where $M=40$, is much smaller than in classical reservoir computing (where its dimension would be approximately $1,000$), making the training of RC less computationally intensive.
@@
Finally, the output matrix $\mathbf{W}_{\mathrm{out}}\in\mathbb{R}^{2\times M}$ maps $\mathbf{\tilde{x}}$ to the target output $\mathbf{y}\in\mathbb{R}^{2}$. Since we do not have access to any values of the committor, these cannot be used as target output for the linear regression. Instead, the only accessible information is whether $\mathbf{x}$ first leads the trajectory to $A$ or to $B$ or if $T_{\mathrm{max}}$ is reached before any of these sets. So the target output is $\mathbf{y}=(y_{1},y_{2})$ where $y_{1}=1$ if $A$ or $T_{\mathrm{max}}$ is reached first and 00 otherwise ; $y_{2}=1$ if $B$ is reached first and 00 otherwise. A Softmax operation then transforms these predictions finally into probabilities, namely: (”probability to reach $A$ or $T_{\mathrm{max}}$ first”, ”probability to reach $B$ first”). The latter corresponds to the committor function.
@@
Because it is a simple linear regression that transforms $\tilde{\mathbf{x}}$ into $\mathbf{y}$, an analytical expression of $\mathbf{W}_{\mathrm{out}}$ can be derived (Lukoševičius and Jaeger, 2009; Gauthier et al., 2021), only depending on the feature vectors $\mathbf{\tilde{X}}$ and associated labels $\mathbf{Y}$:
@@
where $\mathbf{I}$ is the identity matrix and $\gamma$ is the Tikhonov regularization parameter. In our setup, $\mathbf{\tilde{X}}$ depends only on $\gamma=10^{-4}$ and on $p=4$.
Finally, an important advantage of using RC is their ability to be trained online. Here, we would like to use as little data as possible to train the RC, so we use only the data provided by TAMS and retrain the RC at every iteration of the algorithm. Let $\mathbf{\tilde{X}}_{k}$ and $\mathbf{Y}_{k}$ be all the feature vectors and all the labels we used up to iteration $k$. If new data become available, we can add the newly computed feature vectors and labels (say $\mathbf{\tilde{x}}$ and $\mathbf{y}$) as new columns of $\mathbf{\tilde{X}}_{k}$ and $\mathbf{Y}_{k}$. The matrices $\mathbf{Y}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}$ and $\mathbf{\tilde{X}}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}$ can then be easily updated:

### v3

@@ 6 Discussion
This is to our knowledge the first study to couple TAMS to a machine learning technique that estimates the committor function on-the-fly. On the one hand, Lucente et al. (2022) had already coupled Adaptive Multilevel Splitting (an algorithm similar to TAMS) to another machine learning technique, the Analogue Markov Chain, which estimated the committor beforehand, thus creating an additional computational cost. On the other hand, direct committor function estimation using machine learning has already been performed on complex climate data (Miloshevich et al., 2023b), but only for short-term probabilistic forecast.
@@
We showed that the committor function can be learned at no additional cost than needed to compute transition probabilities. We compared our results to a combination of TAMS with a physics-informed score function and showed that transition probabilities and mean first-passage times were consistent. In the case of F-transitions, TAMS-$R$ also yields on average a similar variance in the results as TAMS-$S$, except for very low noise. Regarding S-transitions, both methods remain consistent for transition probabilities within $1000$ years down to $10^{-6}$. Furthermore, TAMS-$R$ consistently converges in fewer iterations than TAMS-$S$, suggesting that, in this case, RCs can learn a more efficient score function than $\Phi_{S}^{S}$. Finally, we showed that the analytical interpretation of the learned committor is consistent in the case of the F-transition with the physics-informed score function.
Concerning the AMOC, we confirmed that the probability of a collapse within a time frame of $100$ years for the F-transition and $1000$ years for the S-transition is not negligible for a large range of model parameters. Moreover, our method yields transition paths that are consistent with those obtained with a physics-informed score function. Finally, some of the terms of the estimated expression of the committor indicate that the RC captures the salt-advection feedback as key in the collapse process.
@@
The next step is to study AMOC tipping in higher-dimensional models like Earth Models of Intermediate Complexity (EMICs) or even Global Climate Models (GCMs). Another rare-event algorithm (Moral and Garnier, 2005; Tailleur and Kurchan, 2006; Lecomte and Tailleur, 2007) has already been applied to EMICs (Ragone et al., 2018; Ragone and Bouchet, 2021; Wouters et al., 2023) and AMOC tipping (Cini et al., 2024). However, it focuses on the sampling of rare events rather than computing transition probabilities. Another interesting data-driven approach is the application of the edge-tracking algorithm to AMOC tipping (Mehling et al., 2024). It finds the critical state through which an AMOC tipping would occur, using a limited number of simulations, and is thus tractable in higher dimensions. This approach provides insight into the dynamics of the system but does not quantify the tipping probability.
@@
Extending our method to complex climate models raises the question of its scaling with the dimension of the model. For TAMS-$R$, this question is twofold: the feasibility of simulating enough trajectories and the scaling of RC itself. Regarding the first point, TAMS is not more computationally demanding than any other rare-event algorithm so their recent successes in EMICs (Cini et al., 2024) are promising for TAMS. When it comes to GCMs (van Westen and Dijkstra, 2023), running TAMS does not seem feasible yet.
@@
However, the current implementation of the RC method poses a scaling problem. Due to the monomials used as basis functions, the size of the feature vector grows exponentially with the number of variables. We can only consider a very limited number of variables, which may be too simple to describe AMOC tipping in a GCM. Miloshevich et al. (2023a) showed that Convolutional Neural Networks (CNN) are suitable for learning the committor from maps of climate variables. However, training CNNs still requires a large amount of data. More generally, this raises the problem of the choice of basis functions for the RC. The same issue arises for another committor estimation method, the Dynamical Galerkin Approximation (Thiede et al., 2019; Finkel et al., 2021; Jacques-Dumas et al., 2023). Strahan et al. (2023) recently used a Feedforward Neural Network to solve the Feynman-Kac formula for the committor. It only requires short trajectories such as those obtained with TAMS but may not solve the issue of the amount of data needed for training.
@@
The interpretability of the RC could also prove useful on the question of dimensionality. We showed in Sect. 5 that in the feature vector of the RC, only a dozen of the terms play a significant role in the estimate. If these can be detected early enough in the learning process, dropping the others would speed up the computation. More generally, neural networks are often considered to be black boxes and difficult to interpret. Here, important variables and their relationship can be highlighted, which may serve as a useful starting point to focus observations or experiments. The efficiency of TAMS may rely on sufficient insight into the model’s dynamics (Lestang et al., 2018), but we show here that an efficient TAMS can also provide such insight.
@@
Finally, the initialization and training process of the RC must be refined. In complex models, transition probabilities may only be computed for a single model setup, which makes transfer learning unusable. The RC could then be initialized from a linear score function that makes it easy to generate surrogate data. Another important issue is the memory of the RC. The current transfer learning process may retain too much information corresponding to very different system parameters since we did not ”clean” its memory during the computations. The same issue may arise if the initialization of the RC is very far from the actual committor. A possible solution is to reinitialize the RC at certain milestones, once a certain number of transitions have been reached, thus using ”better” data.
@@ Acknowledgements.
    Simulate $N$ trajectories $\left(\mathbf{X}^{(1)},\dots,\mathbf{X}^{(N)}\right)$ starting in set $A$ until $T_{\mathrm{max}}$ or until they reach set $B$. Repeat steps $3-8$ until at least $N-n_{c}$ (or $N$ if $n_{c}=1$) trajectories have reached $B$.
@@
    For trajectory $i$, the maximum value of the score function is called $Q_{i}$. Let $I_{k}=\{i\in[1,N]\ |\ Q_{i}\in\{Q^{*}_{1},...,Q^{*}_{n_{c}}\}\}$, where $\{Q^{*}_{j}\}_{j\in[1,n_{c}]}$ are the $n_{c}$ lowest values of $Q_{i}$. Since the system is discrete, $I$ may contain more than $n_{c}$ elements.
@@
    Randomly select a trajectory $\mathbf{X}^{(r)}$ (such that $r\notin I_{k}$). We call $\tau$ the first time so that $\Phi(\mathbf{X}^{(r)}(\tau))\geq Q_{i}$.
@@
    Set $\mathbf{X}^{(i)}([1,\tau])=\mathbf{X}^{(r)}([1,\tau])$.
@@
Let $N_{B}$ the number of trajectories that reached $B$ and $K$ and the number of iterations. The estimated transition probability from $A$ to $B$ before $T_{\mathrm{max}}$ is given by (Rolland and Simonnet, 2015; Rolland, 2022):
$\hat{\alpha}=\frac{N_{B}}{N}w_{K}=\frac{N_{B}}{N}\prod_{k=1}^{K}w_{k}=\frac{N_{B}}{N}\prod_{k=0}^{K-1}\left(1-\frac{\#I_{k}}{N}\right)$ (8)
The optimal variance, obtained when the score function is the committor function reads:
$\sigma^{2}_{\mathrm{id}}=\frac{<\hat{\alpha}>^{2}}{N}\left(<K>\frac{n_{c}}{N-n_{c}}+\frac{N-<N_{B}>}{<N_{B}>}\right)$ (9)
where <>$\cdot$>< ⋅ > represents an average over $M$ runs of TAMS.
@@
Reservoir Computing (Jaeger, 2001; Lukoševičius and Jaeger, 2009) is a kind of neural network designed to represent dynamical systems and temporal processes better than feedforward networks. The idea is to transform a nonlinear problem in the original phase space into a linear problem by embedding input data into a larger-dimensional space where linear regression is performed. Classical reservoir computers are universal approximators (Gonon and Ortega, 2020) but pose certain problems. For instance, their performance can vary depending on the realizations of the reservoir, which is usually a fixed random matrix.
@@
The “Next-Generation” Reservoir Computing (here abbreviated RC) developed by Gauthier et al. (2021) is also an universal approximator (Gonon and Ortega, 2020; Hart et al., 2021) but embeds input data into a large-dimensional space using a family of nonlinear basis functions instead of a random matrix. We provide below only the important steps to build an RC for our problem. Details of the architecture can be found in Gauthier et al. (2021) and in Jacques-Dumas et al. (2023) for the setup and choice of parameters.
@@
The RC receives at each time step a vector $\mathbf{x}\in\mathbb{R}^{4}$ containing $S_{n}$, $S_{ts}$, $D$ and the time $t$. The family of non-linear functions consists of all products of power $p=4$ of these variables. The resulting feature vector $\mathbf{\tilde{x}}\in\mathbb{R}^{d}$, with $d=40$ is much smaller than in classical reservoir computing (where its dimension would be approximately $1,000$), making the training of RC more efficient.
@@
The output matrix $\mathbf{W}_{\mathrm{out}}\in\mathbb{R}^{2\times d}$ maps $\mathbf{\tilde{x}}$ to the target output $\mathbf{y}\in\mathbb{R}^{2}$. Since the committor is inaccessible, its values cannot be used as target for the linear regression. We only know whether $\mathbf{x}$ first leads the trajectory to $A$ or to $B$ or if $T_{\mathrm{max}}$ is reached before any of these sets. The target output is $\mathbf{y}=(y_{1},y_{2})$ where $y_{1}=1$ if $A$ or $T_{\mathrm{max}}$ is reached first and 00 otherwise ; $y_{2}=1$ if $B$ is reached first and 00 otherwise. A Softmax operation then transforms these predictions into probabilities, namely: (”probability to reach $A$ or $T_{\mathrm{max}}$ first”, ”probability to reach $B$ first”). The second is an estimate of the committor.
@@
Since $\tilde{\mathbf{x}}$ and $\mathbf{y}$ are matched through linear regression, an analytical expression of $\mathbf{W}_{\mathrm{out}}$ can be derived (Lukoševičius and Jaeger, 2009; Gauthier et al., 2021), only depending on the feature vectors $\mathbf{\tilde{X}}$ and associated labels $\mathbf{Y}$:
@@
where $\mathbf{I}$ is the identity matrix and $\gamma$ is the Tikhonov regularization parameter.
Finally, an important advantage of using RC is their ability to be trained online, for instance to update $W_{\mathrm{out}}$ at each iteration of TAMS. Let $\mathbf{\tilde{X}}_{k}$ and $\mathbf{Y}_{k}$ be all the feature vectors and all the labels used up to iteration $k$. If new data become available, the newly computed $\mathbf{\tilde{x}}$ and $\mathbf{y}$ become new columns of $\mathbf{\tilde{X}}_{k}$ and $\mathbf{Y}_{k}$. The matrices $\mathbf{Y}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}$ and $\mathbf{\tilde{X}}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}$ are then updated: